OpenStudy (anonymous):

5-x-x= -1 (& how did you get the answer?)

OpenStudy (anonymous):

5-2x=-1 add 1 to both sides 6-2x=0 add 2x to both sides (-2x is -x-x simplified) 6=2x divide by two x=3

OpenStudy (anonymous):

THANK YOU SO MUCH!

OpenStudy (anonymous):

yup no problem